# Training Large Neural Networks With Low-Dimensional Error Feedback

### v3

@@ 4.   Nonlinear networks and complex architectures
To investigate this question, we applied our approach to visual transformer architectures trained for image recognition on the CIFAR-10 dataset. Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer. Feedback matrices were trained using the LDFA framework (see the appendix for details). The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
Our results indicate that, consistent with our findings in both nonlinear and convolutional networks, an error signal with dimensionality as low as the task dimensionality (i.e. the number of output classes) is sufficient to train transformer architectures effectively (Figure 3f). Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$. Notably, training transformers with low-dimensional feedback using LDFA not only matched but, in some cases, consistently outperformed backpropagation by a small margin. This improvement was observed even when hyperparameters were independently optimized for both BP and LDFA. The improved generalization could be understood as a regularization effect achieved by constraining the feedback. However, since we observe it only in the transformer architecture, it is not a general property of a low-dimensional error signal, and we leave further investigation to a future study.
Figure 4: Receptive fields (RFs) are shaped by feedback dimensionality rather than feedforward constraints. (a) Model of the early ventral visual stream. Top: Retinal-to-cortical connectivity is modeled as a bottleneck in the network; adapted from [29]. Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality. Bottom: Using direct LDFA (dLDFA), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck. (b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top. Left: No bottleneck. Right: Bottleneck at the retinal output, which induces center-surround RFs. (c) Training the retinal output with LDFA. Restricting the rank of the feedback matrix produces center-surround RFs. (d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dLDFA. The presence of high-dimensional error signals results in oriented RFs.
@@ 5.   Error dimensionality shapes neural receptive fields
We have shown that neural networks can be efficiently trained using minimal error signals comparable to task dimensionality. Here, we investigate how error dimensionality affects neural representations, providing insights into receptive fields observed in the brain.
Lindsey et al. found that in convolutional models of the visual system, narrow feedforward bottlenecks between the retina and the brain led to center-surround receptive fields in the retinal layer, similar to mammals [29]. Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders. We hypothesize that these effects are due to constraints on the error signal reaching the retina rather than the feedforward bottlenecks themselves.
To test this hypothesis, we trained a model similar to that of Lindsey et al.  [29], but with full-width layers throughout. Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with LDFA (Figure 4c), thereby isolating the impact of error dimensionality on neural representations. We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix). This experiment demonstrates that error dimensionality influences neuronal tuning and neural representations. Specifically, lower-dimensional error signals promote higher symmetries in emergent receptive fields.
Importantly, the architecture studied by Lindsey et al. is not a faithful model of the early visual system. In particular, there is no evidence that error signals propagate from the cortex or thalamus back to the retina. Instead, it is believed that the receptive fields in the retina emerge through unsupervised learning [31]. Nevertheless, our results provide an explanation for the receptive fields observed by Lindsey et al. Furthermore, it may lead to an alternative explanation of learning in the visual pathway that uses simplified signaling (e.g., neuromodulators). Most importantly, our findings show a relation between tuning curves and error pathways.

### v4

@@ 4.3 Transformer-based architectures
To test this, we applied our low-rank feedback scheme to a vision transformer trained on CIFAR-10. All weight matrices in the architecture, including the key, query, and value projections, were trained using low-dimensional error feedback from the subsequent layer (see Appendix for details). For simplicity, we used the same feedback rank for all matrices, so the error dimensionality is controlled solely by the rank $r$ of each learned backward map. Performance as a function of feedback rank is shown in fig.˜3b.
Despite the complexity of self-attention, transformers remained trainable with error dimensionality as low as the task dimensionality. In particular, rank $r=24$ feedback yielded performance on CIFAR-10 comparable to standard backpropagation (Figure 3b,c), consistent with our observations in linear and convolutional networks.
4.4 Scaling and efficiency gains
The results above show that low-dimensional feedback suffices to train deep nonlinear architectures, including convolutional networks and transformers, to near backpropagation-level accuracy. This has two implications: it decouples how performance scales with feedforward width from how the feedback pathway must scale, and it opens a regime where training can be more computationally efficient than standard backpropagation.
Scaling with network width.
Constraining the feedback to be low-rank changes the credit-assignment channel but leaves the feedforward model class unchanged. The representational capacity remains set by the width and depth of the forward network, not by the rank of the feedback matrices. Consistent with this, increasing the size of the feedforward network improves accuracy even when the feedback rank is held fixed (fig.˜3c top). At the same time, our linear analysis above indicate that the rank required to match backpropagation accuracy is controlled primarily by task dimensionality $d$ (e.g., performance degrades when $r<d$). Together, these observations imply that one can benefit from overparameterized feedforward representations while keeping the feedback dimensionality tied to task complexity rather than to network width.
Computational efficiency.
A direct benefit of low-rank feedback is a reduced cost of propagating errors through depth. For a layer with widths $n_{l}$ and $n_{l+1}$, backpropagation multiplies by $W_{l}^{\top}$ at cost $O(n_{l}n_{l+1})$ per example (or per token). With low-rank feedback $B_{l}=Q_{l}P_{l}$ of rank $r$, the same operation factors into $P_{l}\boldsymbol{\delta}{l+1}$ followed by $Q_{l}(\cdot)$, at cost $O(r(n_{l}+n_{l+1}))$. Accordingly, the FLOPs associated with backward error propagation decrease linearly with the effective error dimensionality $r$.
Total training cost reflects three additional effects. First, the factors $Q_{l}$ and $P_{l}$ must be updated, but these updates depend only on $W_{l}$ and do not scale with batch size, so they can be amortized by updating the feedback factors at longer intervals. Second, extremely small ranks can slow convergence; in the regime where final performance matches backpropagation, learning curves are comparable (fig.˜3c top), whereas at very low ranks, additional epochs may be required. Third, the backward pass is only part of training, which also includes the full-rank forward pass and the weight updates.
To capture these effects, we measure the total number of FLOPs required to reach $90\%$ of the final test accuracy, jointly accounting for backward-pass savings, feedback-factor updates, and any change in convergence time. This yields a U-shaped dependence on rank: at large $r$, per-iteration cost approaches backpropagation, whereas at very small $r$ slower convergence dominates. For a vision transformer on CIFAR-10, an intermediate rank achieves the best trade-off, reducing total compute (by about $20\%$ in our experiments) while maintaining backpropagation-level accuracy (fig.˜3c bottom).
Figure 3: Low-dimensional feedback in nonlinear vision architectures. (a) Convolutional networks. A VGG-like CNN trained on CIFAR-10 with learned low-rank feedback achieves near-backpropagation (BP) test accuracy even when the backward/teaching channel is strongly compressed. Bars sphow final accuracy for BP and for low-dimensional feedback with feedback rank set to a fraction of the layer width (r=n/2,n/4,n/8r=n/2,\,n/4,\,n/8, where $n$ denotes the number of channels in the corresponding block). (b) Vision transformers (ViT). Top, schematic of a transformer block and multi-head self-attention with low-rank feedback channels; low rank feedback used to train all weights, including MLP and attention weights. Bottom, training curves (test accuracy versus training step) for BP (black) and low-dimensional feedback with a shared feedback rank $r$ across all linear maps (color-coded); inset shows the corresponding training loss. (c) Scaling and efficiency in ViT. Top, holding the feedback rank fixed ($r=36$) while reducing the embedding dimension decreases accuracy, indicating that performance continues to benefit from larger feedforward representations even under a fixed low-dimensional error channel. Bottom, ViT test accuracy (bars) and total compute required to reach $90\%$ of the final accuracy (line, TFLOPs) as a function of feedback rank, revealing an intermediate-rank regime that preserves BP-level accuracy while reducing end-to-end training compute (about $20\%$ in these experiments).
5.  Local learning rules
The factorized feedback-learning scheme above achieves backpropagation-level performance with an error channel of rank $r\ll n_{l}$, and yields tangible compute savings. However, at its core it remains a surrogate for weight transport since the update of $(Q_{l},P_{l})$ is driven by the corresponding forward matrix $W_{l}$ through the objective $\|Q_{l}P_{l}-W_{l}^{\top}\|_{F}^{2}$. Implementing such “learning through distance” would require synapses in the feedback pathway to access parameters in a separate, remote pathway, which is difficult to reconcile with known biological plasticity mechanisms. Moreover, explicitly approximating $W_{l}^{\top}$ presupposes a parallel forward route whose weights are being transposed; this assumption breaks down when error signals arrive through indirect pathways that do not mirror the feedforward circuit, as in direct feedback alignment [46]. These limitations motivate a complementary perspective: rather than learning a proxy for $W_{l}^{\top}$, we seek a genuinely local mechanism that learns a useful low-dimensional error representation while keeping the remaining feedback circuitry fixed.
Our linear analysis suggests that, under a rank constraint, the crucial degree of freedom is not an element-wise approximation of $W_{l}^{\top}$ but the error subspace that the feedback pathway can represent and transmit. Writing $B_{l}=Q_{l}P_{l}$, the matrix $P_{l}\in\mathbb{R}^{r\times m}$ projects output-space errors into an $r$-dimensional channel, so the columns of $P_{l}^{\top}$ span the subspace of output error directions that can be assembled upstream. When this subspace is poorly aligned with the task, the stationarity condition becomes underdetermined and training can converge to spurious fixed points. Conversely, when $P_{l}^{\top}$ spans task-relevant directions, the feedback pathway preserves informative components of the error while discarding directions that are redundant for learning.
Ideally, $P_{l}$ has $r$ orthonormal rows ($P_{l}P_{l}^{\top}=I_{r}$), so that $P_{l}^{\top}P_{l}$ is the rank-$r$ projector onto the transmitted subspace in the downstream error space. Thus, $P_{l}$ learns the top-$r$ subspace of the incoming teaching signal to layer $l$ (i.e., the vector that $P_{l}$ multiplies). Concretely, we want the transmitted subspace to align with the top principal components of the layerwise error, given by the eigenvectors of $\frac{1}{p}\sum_{\mu}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu\top}$, where $\boldsymbol{\delta}_{l+1}^{\mu}$ denotes the locally available teaching/error signal layer $l+1$ for example $\mu$. Specifically, at the output layer, the error is given by $\boldsymbol{\delta}_{L}^{\mu}=\partial\mathcal{L}^{\mu}/\partial\hat{\boldsymbol{y}}^{\mu}$, while for hidden layers it is the error signal broadcast through the constrained feedback pathway. This choice allocates the limited rank of the feedback channel to the directions along which the network is currently making the biggest mistakes, and it uses a signal that is locally available wherever feedback arrives.
To learn the principal components of the error using online updates, we use a multi-component Oja rule driven by the locally available teaching signal [48],
$\Delta P_{l}^{\mu}=\eta\,P_{l}\,\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu\top}\,(I-P_{l}^{\top}P_{l})-\lambda P_{l},$ (15)
where $I$ is the identity matrix.
Eq. (15) is a matrix-valued Oja update for online principal-subspace learning. A Hebbian term driven by the teaching signal $\boldsymbol{\delta}_{l+1}^{\mu}$ increases the components of $P_{l}$ along directions of high error variance, while the remaining factors stabilize the dynamics and set the overall scale. In particular, the factor $(I-P_{l}^{\top}P_{l})$ promotes row-orthonormality of $P_{l}$, and the decay term $\lambda P_{l}$ controls the magnitude of the weights.
Although the orthonormalizing factor is written compactly as a matrix product, its role is to express within-population competition and normalization in the $r$-dimensional error channel. In Oja’s original single-component rule, the stabilizing term can be interpreted as a postsynaptic-activity–dependent normalization of an otherwise Hebbian increment, yielding an online update that converges to the leading principal component [48]. Extending PCA learning to multiple components necessarily requires a mechanism that prevents different components from collapsing onto the same direction, which is achieved by introducing competition among the units representing the learned subspace.
Several biologically motivated constructions realize this competition using only local interactions within the postsynaptic population [58, 29, 30, 44, 27, 51]. In particular, Sanger’s generalized Hebbian algorithm implements multi-component PCA by combining Hebbian plasticity with within-layer interactions that subtract already-learned projections, thereby enforcing orthogonality through local competition among units [58]. More generally, similarity-matching formulations derive networks in which feedforward synapses learn Hebbianly while lateral synapses learn anti-Hebbianly, and the lateral dynamics locally decorrelate and normalize neural activities, effectively playing the role of the orthonormalization implicit in (15[51, 52]. In this sense, (15) should be read as a compact description of a circuit-level computation in which synaptic updates depend on pre- and postsynaptic activity, while orthogonality and normalization are maintained by local competition within the error-channel population.
Finally, Because $P_{l}$ adapts gradually, it tracks the dominant modes of the nonstationary teaching signal; the forward weights then reduce those modes, causing the subspace to rotate as learning progresses.
Together with the layer-wise weight update (13), this yields a local learning scheme in which feedback synapses update from signals available at their pre- and postsynaptic terminals, and stabilization is mediated by within-layer normalization and competition rather than by nonlocal parameter access.
We refer to this scheme as low-dimensional feedback alignment (LDFA). In what follows, we show that learning $P_{l}$ alone, with $Q_{l}$ fixed and random, is sufficient to reach backpropagation-level accuracy; as in Feedback Alignment [38], the forward weights can then align to the fixed mixing induced by $Q_{l}$.
5.1 Linear analysis
We analyze the coupled learning dynamics in the one-hidden-layer linear network of Section 3, using the same rotation into the singular basis of the input-output correlation $\Sigma_{io}=USV^{\top}$. Writing $W_{1}=\bar{W}_{1}V^{\top}$, $W_{2}=U\bar{W}_{2}$, and $P=\bar{P}U^{\top}$ yields an effective feedback $B=QP=Q\bar{P}U^{\top}$, with $Q$ fixed and $\bar{P}(t)$ learned. In the continuous-time limit, the forward dynamics in the rotated basis are exactly those of eq.˜5, with the sole modification that $\bar{B}(t)=Q\bar{P}(t)$ is now time-dependent.
The new ingredient is the evolution of $\bar{P}$ under the error-driven Oja rule (15). For squared loss and whitened inputs, the output-layer error in the rotated coordinates satisfies $\bar{\boldsymbol{\delta}}^{\mu}(t)=\bar{E}(t)\,\bar{\boldsymbol{x}}^{\mu}$, where $\bar{\boldsymbol{x}}^{\mu}=V^{\top}\boldsymbol{x}^{\mu}$ and $\bar{E}(t)=S-\bar{W}_{2}(t)\bar{W}_{1}(t)$ is the residual in the singular basis. Averaging over samples, therefore, gives an error covariance $\bar{E}(t)\bar{E}(t)^{\top}$, and the continuous Oja update can be written as
$\tau\frac{d\bar{P}}{dt}=\bar{P}\,\bar{E}\bar{E}^{\top}\,(I-\bar{P}^{\top}\bar{P})-\lambda\bar{P}.$ (16)
Equations (5) and (16) thus define a coupled first-order dynamical system for $(\bar{W}_{1},\bar{W}_{2},\bar{P})$.
This system has a straightforward interpretation. The Oja flow (16) steers the transmitted subspace toward the leading eigenvectors of the current error covariance, so the rank-$r$ channel allocates its constrained capacity to the modes that dominate the residual. As learning progresses, this mechanism tracks the dominant modes of the current residual, which in this solvable setting biases learning toward the leading singular directions of $\Sigma_{io}$ (Figure 4a).
Figure 4: Low-dimensional feedback alignment (LDFA) as a synaptically local error pathway. (a) Linear-network analysis ($n=128$, $k=64$, $m=64$) with factorized feedback $B=QP$. Solid curves show simulations and dashed curves show the analytical prediction for the evolution of mode-wise singular values. With full-rank feedback ($r=m$), all task modes are recovered, whereas with rank-constrained feedback ($r<m$) learning prioritizes the dominant task modes selected by the Oja dynamics. (b) Multilayer perceptrons on CIFAR. Top: separating feedforward capacity from feedback dimensionality. A wide MLP reaches high accuracy even when the teaching signal is strongly compressed (LDFA with low rank $r$), whereas a much narrower MLP performs substantially worse despite having an unconstrained (full-rank) error pathway (BP). Bottom: varying task output dimensionality by subsampling classes from CIFAR-100 shows that the minimal rank required to match BP scales primarily with task dimension (number of classes), not with hidden-layer width (all layers use the same rank $r$). (c) Direct LDFA implements direct error projection by broadcasting a low-dimensional teaching signal to each layer either from the output layer (solid) or the penultimate layer (dashed), approaching BP-level accuracy as rank increases. (d) Convolutional networks trained with LDFA. A 4-block VGG-style CNN attains BP-level performance with low backward rank even when the widest blocks contain 512 channels. Inset: constraining all layers to a fixed fraction of their width ($r=n/2,n/4,n/8$) yields graceful degradation.
5.2 Deep nonlinear networks
We next test whether the local rule extends beyond the linear setting to deep nonlinear networks. Following the same procedure in section˜4, we train MLPs on CIFAR using the same constrained feedback architecture as above, where each layer receives an $r$-dimensional error signal through a fixed random mixing $Q_{l}$, while the projection $P_{l}$ is learned locally by the error-driven Oja update (15). Across architectures, we find that as with the normative approach in section˜3, rank required to match backpropagation tracks task dimensionality rather than network size. Concretely, when we subsample CIFAR-100 to classification problems with 50, 65 or 100 classes and constrain all feedback matrices to rank $r$, performance approaches full backpropagation once $r$ reaches the output dimension defined by the number of classes (Fig. 4b).
5.3 Low-dimensional error broadcasting
A key limitation of the normative feedback-learning rule in section˜3 is its explicit dependence on the forward weights $W_{l}$, in the minimization of eq.˜8. This effectively assumes reciprocal connections between consecutive layers, and therefore cannot be applied to error pathways that bypass intermediate areas. In biological circuits, however, instructive signals are known to broadcast through indirect and long-range projections, while a mechanism that rely on local reciprocals has yet been found. A canonical example is the dopaminergic reward-prediction-error signal, which is delivered to many downstream targets via widespread projections [45, 62, 65].
Direct feedback alignment (DFA) was introduced precisely to relax the requirement of reciprocal connectivity by broadcasting the output error directly to each hidden layer through fixed random feedback matrices [46]. In its simplest form, each layer receives a teaching signal computed in parallel from the output-layer error, enabling updates without sequential backward transport through intermediate layers. While DFA can train deep networks in some regimes, its naive formulation typically exhibits a substantial performance gap relative to backpropagation on modern vision benchmarks, and convolutional architectures are a particularly challenging case [4, 14].
Our local rule removes the architectural constraint that prevented the normative approach from operating in this regime. Concretely, for each layer $l$ we broadcast the error from a chosen source layer $l^{\prime}$ (often $L$) directly to layer $l$ using a low-rank matrix,
$\boldsymbol{\delta}_{l}=\big(Q_{ll^{\prime}}P_{ll^{\prime}}\boldsymbol{\delta}_{l^{\prime}}\big)\odot f^{\prime}(a_{l}),\qquad l<l^{\prime}\leq L$ (17)
where $l^{\prime}$ is an efferent layer to $l$ or the output error $\delta_{L}$. Same as above, $Q_{ll^{\prime}}$ is fixed and random and $P_{ll^{\prime}}\in\mathbb{R}^{r\times n_{l^{\prime}}}$ is learned by the local Oja update eq.˜15 (with $\boldsymbol{\delta}_{l^{\prime}}$ as its driving signal). This construction preserves the defining feature of DFA, namely parallel error broadcast through non-reciprocal pathways, while endowing the broadcast channel with an adaptive, low-rank projection that allocates its limited capacity to the dominant directions of the task error. Because the update of $P_{ll^{\prime}}$ is unsupervised and depends only on the locally available teaching signal, it applies unchanged when the feedback edges skip layers or originate from alternative sources (e.g., the penultimate layer), and it does not require access to the forward weights.
Empirically, this direct error projection closes much of the gap between DFA and backpropagation in deep nonlinear networks. In particular, learning $P_{l}$ with $Q_{l}$ fixed suffices to train deep MLPs and CNNs with direct error broadcasts to backpropagation-level performance (Fig. 4c), including regimes in which standard DFA is known to struggle. This demonstrates that the local subspace-learning mechanism is compatible with non-mirrored, indirect error pathways, strengthening the biological motivation of the framework while retaining the benefits of low-dimensional feedback.
5.4 Convolutional networks
A particularly stringent test is the convolutional setting. In deep CNNs, naive alignment schemes often exhibit a substantial gap relative to backpropagation unless augmented with additional mechanisms that improve alignment or stabilize training. Our low-dimensional feedback framework remains effective in this regime: we train CNNs reliably and reach backpropagation-level accuracy at comparable rank (Fig. 3b), despite the combined challenges of deep credit assignment, weight sharing, and a restricted error channel.
In all CNN experiments we learn the error projection $P_{l}$ using the local Oja update (15), which is the core mechanism that allocates the rank-$r$ channel to task-relevant error directions. Convolutional architectures, however, are a particularly stringent regime for alignment-based learning, and prior work on feedback alignment has typically found that allowing some adaptation of the feedback weights substantially improves robustness and accuracy in deep CNNs [34, 1]. Following this established practice, we therefore also allow the mixing matrix $Q_{l}$ to adapt via a simple synaptically local Hebbian rule,
$\Delta Q_{l}^{\mu}=\eta\,\boldsymbol{h}_{l}^{\mu}\big(P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\big)^{\top}-\lambda Q_{l},$ (18)
which preserves the low-rank structure and does not introduce weight transport or any dependence on forward parameters. Importantly, (18) remains local: the presynaptic activity $\boldsymbol{h}_{l}^{\mu}$ and the postsynaptic teaching signal $P_{l}\boldsymbol{\delta}_{l+1}^{\mu}$ are available at the feedback synapse, and the update depends only on these signals and the synaptic weight. Empirically, this augmentation is consistent with the FA literature on challenging vision models and improves performance in convolutional architectures while retaining the central feature of LDFA: learning a low-dimensional error subspace through $P_{l}$.
Figure 4d shows that training CNNs with low-dimensional error and local-only learning rules can match backpropagation performance, while maintaining more biological integrity.
@@ 6.  Error dimensionality shapes neural receptive fields
Figure 5: Receptive fields are shaped by feedback dimensionality. (a) Simplified model of the early ventral visual stream (retina $\rightarrow$ VVS) adapted from [40], and two manipulations that decouple feedforward bottlenecks from the teaching pathway. Top: an anatomical bottleneck limits the retinal output. Middle: the feedforward architecture is left full-width, but the teaching signal to the retinal layer is restricted by a low-rank feedback map (LDFA). Bottom: the feedforward bottleneck is reinstated, while the retinal layer receives a high-dimensional teaching signal via direct error projection (dLDFA; Eq. (16)), bypassing backward compression. (b) Backpropagation reproduces the Lindsey et al. phenomenology: removing the retinal bottleneck yields oriented, edge-like receptive fields, whereas a narrow retinal bottleneck yields center-surround receptive fields. (c) With the feedforward pathway unconstrained but feedback to the retina rank-limited, LDFA reliably produces center-surround receptive fields; decreasing the feedback rank $r$ (examples $r=32,4,2$) increases their effective rotational symmetry. (d) Conversely, with a feedforward retinal bottleneck but high-dimensional error delivered directly to the retina using direct LDFA (dLDFA), receptive fields become oriented. Together, these manipulations show that the dimensionality and routing of the teaching signal can reproduce, and even override, receptive-field structure previously attributed to feedforward anatomical constraints.
At the heart of our framework is the ability to decouple the error-feedback pathway from the feedforward pass, and to control the feedback channels independently. In Secs. 3 and 4 we used this decoupling to study how restricting the backward pathway controls the efficiency of learning, and in Sec. 5 we derived local update rules (including convolutional networks) that implement these constraints using synapse-local information. Here we ask a complementary question: beyond efficiency, does the structure of the error pathway shape the representations learned in early layers? Concretely, we study how receptive fields of shallow layers change when we constrain the dimensionality of the error signal reaching them.
The relation between architecture and learned receptive fields has been studied before. Lindsey et al. showed that in anatomically constrained convolutional models, narrowing the retinal output bottleneck yields center-surround receptive fields in the retinal layer, whereas widening it yields orientation-selective receptive fields [40]. We hypothesize that this transition is driven primarily by the dimensionality of the teaching signal available to the retinal layer. With constrained feedback channels, the compressed error signal biases learning toward higher-symmetry solutions, even when the feedforward pathway is left unconstrained.
LDFA provides a direct way to test this hypothesis because it lets us constrain the backward pathway while keeping the forward architecture fixed, and, via direct error projection (Eq. (17)), to manipulate error delivery independently of feedforward bottlenecks. We first remove the feedforward bottleneck by using full-width layers throughout, and restrict only the backward pathway by training a rank-$r$ feedback matrix to the retinal layer using LDFA (Fig. 5a). We then visualize the resulting retinal receptive fields using standard feature-visualization methods [20] (Fig. 5b,c). Despite the unconstrained feedforward architecture, restricting the feedback rank reliably produces center-surround receptive fields, and decreasing $r$ increases their effective rotational symmetry.
We next perform the complementary manipulation. We reinstate a narrow feedforward bottleneck as in [40], but ensure that the retinal layer receives a high-dimensional error signal by using direct error projections (dLDFA; Eq. (17)), thereby bypassing any backward compression (Fig. 5a bottom). In this setting, the retinal receptive fields become orientation selective (Fig. 5d), matching the wide-bottleneck regime of [40]. Together, these experiments show that the feedback pathway can reproduce, and even override, receptive-field phenomenology previously attributed to feedforward anatomical constraints and that error dimensionality acts as a powerful inductive bias on early-layer selectivity.
The Lindsey et al. architecture is intentionally simplified and should not be read as a mechanistic model of retinal development; in particular, there is no evidence for supervised error signals propagating from cortex or thalamus back to the retina, and early retinal representations are widely thought to be shaped largely by local and unsupervised mechanisms [72]. Our claim is therefore methodological and conceptual. Within a framework in which forward and backward passes are decoupled and implementable with local rules, the dimensionality and routing of learning signals can act as an inductive bias on receptive-field structure. This perspective applies most directly to early sensory stages that plausibly receive low-dimensional instructive signals (e.g., via neuromodulation or top-down pathways), and suggests that constraints on the error pathway may shape emergent selectivity at least as strongly as constraints on the feedforward circuit itself. This complements goal-driven modeling of sensory cortex by highlighting the backward pathway as an additional axis of inductive bias, alongside architecture and objective [70].